Consider the simple moving average model Xt = 0.02 + Wt − 0.4Wt−1, where Wt is a sequence of i.i.d. normal random variables with mean zero and variance 4. What is the mean of Xt? What is the variance of Xt. Show working
Answer #1

Consider the simple moving average model Xt = 0.02 + Wt − 0.4Wt−1, where Wt is a sequence of i.i.d. normal random variables with mean zero and variance 4. What is the mean of Xt? What is the variance of Xt. Show working

Xt = 0.02 + Wt − 0.4Wt−1

Wt ~ N(0,4)

mean of Xt= E(Xt)

=E( 0.02 + Wt − 0.4Wt−1)

=E(0.02)+E(Wt)-0.4*E(Wt-1)

=0.02+0-0.4*0 since E(constant)=constant

=0.02

mean of Xt=0.02

Variance of Xt= V(Xt)

=V( 0.02 + Wt − 0.4Wt−1)

=V(0.02)+V(Wt)-V(0.4*Wt-1)

=0+4+0.4^2*4 since Var(constant)=0

=4+0.64

=4.64

Answer:

mean of Xt=0.02

Variance of Xt = 4.64